OpenStudy (anonymous):

Which of the following equations have the same solution? Solve all three equations separately and explain each step. I. 2x + 4 = 3x – 11 II. 3x + 3 = 51 III. the quantity of 2x plus 1, divided by 3 equals the quantity of x plus 6, divided by 2

OpenStudy (texaschic101):

1. 2x + 4 = 3x - 11 -- subtract 3x from both sides 2x - 3x + 4 = -11 --- subtract 4 from each side 2x - 3x = -11 - 4 -- combine like terms -x = - 15 x = 15 2. 3x + 3 = 51 -- subtract 3 from both sides 3x = 51 - 3 3x = 48 -- divide both sides by 3 x = 48/3 x = 16 3. (2x + 1) /3 = (x + 6)/2 cross multiply 3(x + 6) = 2(2x + 1) -- distribute 3x + 18 = 4x + 2 -- subtract 4x from both sides 3x - 4x + 18 = 2 -- subtract 18 from both sides 3x - 4x = 2 - 18 -- combine like terms -x = -16 x = 16